OpenStudy (anonymous):

How do I solve 4^x = 1/2 ^x-3 a.5 b.9 c.12 d.1

jimthompson5910 (jim_thompson5910):

Hint: \[\Large 4^{x} = \left(\frac{1}{2}\right)^{x-3}\] \[\Large (2^2)^{x} = \left(2^{-1}\right)^{x-3}\] \[\Large 2^{2x} = 2^{-1(x-3)}\] I'll let you finish up.